Ernie's Electrical is evaluating a project which will increase sales by $50,000 and costs by $30,000. The project will cost $150,000 and will be depreciated straight-line to a zero book value over the 10 year life of the project. The applicable tax rate is 34%. What is the operating cash flow for this project in year 1?
Answer #1
Hi, If you like my answer, please rate my answer first and according to my answer...that way only I can earn points. Thanks

Net Income each year = (50000 - 30000 - 150000/10) * (1- 34%) = $3300
So, after tax cash flow = $3300 + $150000/10 = $18300

So operating cash flow = $18300